OpenStudy (lgbasallote):

how do you check if the general solution obtained in a differential equation is correct?

OpenStudy (anonymous):

Plug and Play

OpenStudy (anonymous):

You should be able to take the derivative as many times as you need to in order to get the original diff eq

OpenStudy (unklerhaukus):

Differentiate the solution you should get back the differential equation

OpenStudy (lgbasallote):

really? that means the DE is just integration??

OpenStudy (anonymous):

That's a way of thinking about it, except that the differential equation is an equation that relates a variable to its derivative. It's kind of the same, but the statement isn't 100% accurate.

OpenStudy (lgbasallote):

why not?

OpenStudy (lgbasallote):

WAHHHH pdfs =_=

OpenStudy (lgbasallote):

well anyway i got what i was asking so thanks lol =))

OpenStudy (unklerhaukus):

you can't open it?

OpenStudy (unklerhaukus):

For this Differential equation\[\frac{\text d y}{\text d x}+\frac{x+y}{x}=0\] I solved and got\[x^2+2xy = k^2 \] Differentiating \[\frac{\text d }{ \text dx}\left(x^2+2xy \right)=\frac{\text d }{ \text dx}\left(k^2\right)\] Returns the original equation \[\frac{\text d y}{\text d x}+\frac{x+y}{x}=0\]